If the stock of money in a given economy is 350,000; long run real output equals 420,000units while the fraction of National Income held by the public is 2/3, what will be the long run price level in the economy?
1
Expert's answer
2019-06-06T08:55:56-0400

M = 350,000, Q = 420,000 units.

If the fraction of National Income held by the public is 2/3, then M = 2/3PQ, so:

350,000 = 2/3×P×420,000,

P = 350,000/280,000 = $1.25.